Find a ratio for the following, and write it as a ratio of integers in lowest terms.
Among 100 coffee drinkers, 36 said that they preferred their coffee black and the rest did not prefer their coffee black. Find the ratio of those who prefer black coffee to those who prefer nonblack coffee.

Answer by stanbon(75887)

Find a ratio for the following, and write it as a ratio of integers in lowest terms.
Among 100 coffee drinkers, 36 said that they preferred their coffee black and the rest did not prefer their coffee black. Find the ratio of those who prefer black coffee to those who prefer nonblack coffee.
--------
36 prefer black coffee
100-36 = 64 prefer nonblack coffee
--------
Ratio of prefer black to prefer nonblack = 36/64 = 9/16
==============
Cheers,
Stan H.